OpenStudy (anonymous):

Simplify: (3a – 4b)(a + b) A. 3a2– ab – 4b2 B. 3a2 – 7ab – 4b2 C. 3a2 – 12ab – 4b2 D. 3a2 + ab – 4b2

OpenStudy (anonymous):

(3a−4b)(a+b) =(3a+(−4b))(a+b) =(3a)(a)+(3a)(b)+(−4b)(a)+(−4b)(b) =3a2+3ab−4ab−4b2 =3a2−ab−4b2

OpenStudy (whpalmer4):

multiply each part of the first quantity (3a-4b) by each part of the second quantity (a+b) collect like terms to simplify.

OpenStudy (anonymous):

So the answer is A. :) I broke it down for you.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem @kira12341234 !! :D

OpenStudy (anonymous):

wow thanks everyone! and @mathpapa thanks for the link!

OpenStudy (anonymous):

lol, you are not the real math papa